### v1

# M-LRM: Multi-view Large Reconstruction Model

Abstract
Despite recent advancements in the Large Reconstruction Model (LRM) demonstrating impressive results, when extending its input from single image to multiple images, it exhibits inefficiencies, subpar geometric and texture quality, as well as slower convergence speed than expected. It is attributed to that, LRM formulates 3D reconstruction as a naive images-to-3D translation problem, ignoring the strong 3D coherence among the input images. In this paper, we propose a Multi-view Large Reconstruction Model (M-LRM) designed to efficiently reconstruct high-quality 3D shapes from multi-views in a 3D-aware manner. Specifically, we introduce a multi-view consistent cross-attention scheme to enable M-LRM to accurately query information from the input images. Moreover, we employ the 3D priors of the input multi-view images to initialize the tri-plane tokens. Compared to LRM, the proposed M-LRM can produce a tri-plane NeRF with $128\times 128$ resolution and generate 3D shapes of high fidelity. Experimental studies demonstrate that our model achieves a significant performance gain and faster training convergence than LRM. Project page: https://murphylmf.github.io/M-LRM/
Figure 1: Our M-LRM can generate novel views of high fidelity and meshes of high quality given single or multiple views as input.
footnotetext: Corresponding authors.
1 Introduction
3D reconstruction from multi-view images stands as a fundamental and critical task that has garnered decades of study, playing a pivotal role in various downstream applications including virtual reality, 3D content generation, and robotic grasping. Achieving complete and high-quality reconstructions typically necessitates capturing images from dense viewpoints that offer comprehensive coverage of the target object. However, traditional approaches relying on dense viewpoint images face significant degradation when confronted with limited inputs, such as few or even single images. This limitation results in distorted and incomplete reconstructions due to the inadequacy of information provided. Despite attempts to address 3D reconstruction from a few images or even a single image, existing methods often suffer from inefficiency or yield low-quality results.
Recently, the introduction of the Large Reconstruction Model (LRM) and its multi-view variant, Instant3D, has demonstrated remarkable results, indicating the feasibility of learning a generic 3D prior for reconstructing objects from single images or limited image sets. However, LRM approaches the task of 3D reconstruction from images by simplistically treating it as an image-to-3D regression problem. This involves directly training a highly scalable transformer on extensive data to learn the image-to-3D mapping. Regardless of whether a single image or multiple images are provided as input, LRM first unpacks feature maps of the input images into tokens and then feeds the tokens into the transformer to inherently learn 3D priors. Consequently, when extending the input from a single image to multiple images, the LRM model fails to exhibit significant performance improvements since it disregards the 3D coherence among the input images.
In this paper, we introduce a novel architecture designed for reconstructing 3D shapes from multiple images in a 3D-aware manner, termed as the Multi-view Large Reconstruction Model (M-LRM). In contrast to the LRM, which primarily relies on regressing 3D shapes from images without considering their interrelations, our model draws inspiration from traditional Multi-view Stereo (MVS) techniques and explicitly learns correspondences among input images to infer the 3D geometry. Our model builds upon the triplane-based transformer architecture introduced by LRM, leveraging its scalability on extensive datasets. Specifically, we propose a new multi-view consistent cross-attention scheme, which empowers M-LRM to infer 3D geometry by meticulously examining semantic consistency across input images. Furthermore, we introduce a geometry-aware positional embedding approach that directly incorporates the 3D information encoded in multiple images into the transformer architecture. These two innovations enable M-LRM to learn 3D representations from images in a 3D-aware manner, thereby producing high-quality geometries.
Extensive experiments demonstrate that our M-LRM can generate high-quality 3D objects. It is capable of producing accurate views and geometry. Additionally, our experiments show that incorporating explicit 3D priors into the LRM model significantly speeds up convergence. In summary, our key contributions are as follows:
  • Our M-LRM addresses the shortcomings of instant3D, which lacks modeling of spatial correlation among multiple input views. This leads to slow convergence and sub-optimal generation results. Instead, our M-LRM leverages the spatial coherence of input images to efficiently generate high-fidelity results.
  • We propose a novel Geometry-aware Positional Embedding mechanism to fully utilize image prior. This helps the randomly initialized tri-plane capture spatial and image priors, enabling faster convergence of our M-LRM.
  • We introduce a multi-view consistent cross-attention scheme to accurately query information from relevant image regions. This incorporates explicit spatial priors, making the 3D generative model geometry aware. As a result, our M-LRM can generate consistent novel views and high-quality geometry.
  • Through extensive experiments, we demonstrate that our M-LRM outperforms all baseline methods in sparse-view reconstruction tasks. It can also generate high-quality 3D objects with a single image input when equipped with a multi-view generation model like Zero123++ shi2023zero123++ .
2 Related Work
Figure 2: Visualization of intermediate results during training process. Compared with existing LRM-based approaches, like Instant3D li2023instant3d , our proposed GCA and GaPE remarkably reduce the convergence time and improve the reconstruction quality by a large margin.
3D Generation. Early methods on image-to-3D mainly employ Generative Adversarial Networks (GANs) to produce 3D representations, like point cloud li2018point , signed distance function (SDF) shi2022deep ; zou2023triplane ; chan2022efficient or neural radiance field (NeRF) gu2021stylenerf ; niemeyer2021giraffe . By incorporating 2D GAN architectures brock2018large ; creswell2018generative ; kang2023scaling ; karras2017progressive ; karras2021alias ; karras2019style ; karras2020analyzing ; zhu2023exploring , 3D GANs have made remarkable advancements to generate 3D-aware content in a feed-forward manner. As a representative attempt, EG3D chan2022efficient employs GANs to generate samples in latent space and introduce a compact triplane representation for rendering novel views of 3D objects. Another research branch in 3D generation utilizes diffusion models (DMs) tewari2024diffusion ; ho2020denoising ; rombach2022high ; song2020score ; po2023state ; ddim ; ddpm ; ldm , which have emerged with unprecedented success. One approach involves training 3D DMs directly using 3D supervision gupta20233dgen ; jun2023shap ; nichol2022point ; ntavelis2023autodecoding ; shue20233d or 2D supervision anciukevivcius2023renderdiffusion ; chen2023single ; gu2023nerfdiff ; karnewar2023holodiffusion ; liu2023zero ; shen2023gina . However, these approaches suffer from 3D inconsistency and poor quality due to the lack of 3D priors. DreamFusion poole2022dreamfusion introduces the Score Distillation Sampling (SDS) loss and generates NeRF under the supervision of pretrained large-scale 2D DMs. Variants of DreamFusion chen2023fantasia3d ; chung2023luciddreamer ; hertz2023delta ; liang2023luciddreamer ; lin2023magic3d ; poole2022dreamfusion ; shi2023mvdream ; tang2023dreamgaussian ; wang2023score ; wang2024prolificdreamer aim to reduce Janus (multi-faced) artifacts and improve geometry quality. However, the prolonged optimization process limits its practical application.
Sparse-view Reconstruction Multiview images provide stereo cues facilitating more accurate and detailed shape generation. Consequently, sparse view-based neural reconstruction methods chen2021mvsnerf ; jain2021putting ; lin2023vision ; long2022sparseneus ; wang2021ibrnet ; yu2021pixelnerf achieve significant improvements in generalization to unseen scenes. However, challenges remain in capturing patterns within large-scale datasets. Scaling up the model and datasets is useful, but it also introduces new challenges, such as increased computational complexity. To address these challenges, recent methods leverage large-scale 3D datasets and well-designed multiview diffusion models (DMs) to generate multiple novel views from a single image shi2023mvdream ; wang2023imagedream ; long2023wonder3d ; liu2023syncdreamer ; tang2023MVDiffusion ; tang2024mvdiffusion++ ; liu2023zero123 ; shi2023zero123++ ; one2345 ; li2024era3d . Subsequent steps involve neural reconstruction and Gaussian kerbl20233d or reconstruction tang2024lgm ; xu2024grm . However, these explicit reconstruction methods face challenges due to multi-view inconsistency and low image resolution. Inspired by hong2023lrm ; wang2023pf ; zou2023triplane , we employ the recently proposed transformed-based large reconstruction model for implicit 3D shape generation.
Large Reconstruction Model The extensive 3D datasets deitke2023objaverse ; deitke2024objaverse have opened up new possibilities for training reconstruction models that can create 3D models from 2D images. LRM hong2023lrm first showcased the potential of using the transformer backbone to map image tokens to implicit 3D triplanes with multi-view supervision, resulting in highly generalizable models. Instant3D li2023instant3d achieves highly generalizable and high-quality single-image to 3D generation by extending LRM to a sparse-view reconstruction model and utilizing the prior in 2D diffusion models. While LGM tang2024lgm and GRM xu2024grm models have replaced the triplane NeRF representation with 3D Gaussians kerbl20233d to improve rendering efficiency, Gaussians have limitations in explicit geometry modeling and high-quality surface extraction. Concurrent works $\text{MVD}^{2}$ zheng2024mvd2 , CRM wang2024crm , and InstantMesh xu2024instantmesh use the same transformer backbone and focus on optimizing mesh representation for high-quality geometry and texture modeling. In contrast, our work focuses on incorporating 3D geometry prior to the memory-consuming transformer attention for efficient training. Moreover, our proposed M-LRM, featuring a novel 3D-aware transformer block, allows significant generalization in 3D shape modeling.
3 M-LRM
Figure 3: Overview of M-LRM. M-LRM is a fully differentiable transformer-based framework, featuring a feature encoder, geometry-aware position encoding and a multi-view cross attention block. Given multi-view images with corresponding camera poses, M-LRM incorporates the 2D and 3D features to efficiently conduct 3D-aware multi-view attention. The proposed geometry-aware position encoding allows more detailed and realistic 3D generation.
Fig. 3 illustrates the overall framework of our M-LRM. Unlike LRM which treats single-image-to-3D as a pure image translation problem, our method fully leverages the geometry prior to link 2D image and 3D triplane, thus achieving faster convergence and being able to generate results of better quality. Specifically, we achieve such advances in two steps. (1) Given multiple views of an object, we first extract the image features with an image encoder and propose Geometry-aware Positional Embeddings to initialize the learnable triplane tokens as detailed in Sec. 3.1. (2) With the learnable triplane tokens, we further adopt Geometry-aware Cross Attention to achieve the image feature injection in a geometry-aware manner as described in Sec. 3.2. Finally, we utilize the NeRF mildenhall2021nerf renderer to generate 3D contents with the generated triplane token.
3.1 Geometry-aware Positional Embeddings
In LRM hong2023lrm and its variants wang2023pflrm ; li2023instant3d , the triplane tokens are randomly initialized without any priors and then optimized in the training. For convenience, we denote the randomly initialized triplane tokens as $\mathbf{T}_{r}$. In this section, we propose Geometry-aware Positional Embedding (GaPE) to better utilize image prior instead of the randomly initialized triplane tokens.
Geometry Volume Construction. Inspired by the previous multi-view reconstruction method long2022sparseneus , we aim to directly learn a geometry volume from the input images which encodes local geometric information based on the hints of multi-view consistency of the images. To cope with this, we first construct a grid volume whose vertices $\mathbf{X}=\{\mathbf{x}^{j}\}_{j\in\mathcal{J}}$ are uniformly sampled in the region of interest to be reconstructed. Considering the $i$-th view among the $N$ input views, with its intrinsic $K_{i}\in\mathbb{R}^{3\times 3}$ and extrinsic $E_{i}\in\mathbb{R}^{3\times 4}$, we back project the vertices to the feature map of the $i$-th view $F_{i}$, and derive the corresponding feature grid ${G}_{i}=\{F_{i}(\pi_{i}(\mathbf{x}^{j}))\}_{j\in\mathcal{J}}$ for the $i$-th view, where $\pi_{i}$ indicates the projection that projecting a 3D vertex onto the $i$-th image plane. The projection $\pi_{i}$ is defined as
$\tilde{\mathbf{u}}^{j}_{i}=\pi_{i}\left(\mathbf{x}^{j}\right)=K_{i}E_{i}\tilde{\mathbf{x}}^{j},$ (1)
where $\mathbf{x}^{j}\in\mathbb{R}^{3\times 1}$ is the 3D coordinate of the grid vertex and $\tilde{\mathbf{x}}^{j}\in\mathbb{R}^{3\times 1}$ is the homogeneous coordinate of $\mathbf{x}^{j}$. $\mathbf{u}^{j}_{i}\in\mathbb{R}^{2\times 1}$ is the 2D coordinate of 2D image space and $\tilde{\mathbf{u}}^{j}_{i}\in\mathbb{R}^{3\times 1}$ is the homogeneous coordinate of $\mathbf{u}_{i}^{j}$. With the obtained 2D coordinates, we could use $\operatorname{grid\_sample}$ to extract corresponding features from the 2D feature map $F_{i}$. After back-projecting all the grid vertices, we obtain $N$ feature grids $\{G_{i}\}_{i=1,\cdot\cdot\cdot,N}$, which are fused together into one unified feature volume $\mathbf{G}$ with a 3D convolution network.
Geometric Triplane Token. The feature volume $\mathbf{G}$ contains the enriched prior produced by the image encoder. To introduce such prior to the triplane tokens, for each axis, we collect all the features of the volume along the axis, where the features of different vertices are concatenated, and obtain three feature maps $\mathbf{g}^{xy},\mathbf{g}^{yz},\mathbf{g}^{xz}$. The feature maps are further fed into three unique 2D convolutional networks to produce the positional embedding $\mathbf{T}_{g}^{xy},\mathbf{T}_{g}^{yz},\mathbf{T}_{g}^{xz}$. Finally, fusing the positional embedding and randomly initialized triplane tokens $\mathbf{T}_{r}$ together yields the final geometry-aware triplane tokens as
$\mathbf{T}_{0}=\mathbf{T}_{g}+\mathbf{T}_{r}.$ (2)
Since the tokens explicitly incorporate the geometric priors of the geometry volume, the tokens are geometry-aware and lead to faster convergence.
3.2 Efficient Geometry-aware Cross Attention
Figure 4: The illustration of Geometry-aware Cross Attention.
In addition to the geometry-aware triplane tokens, we introduce a novel cross-attention mechanism, named Geometry-aware Cross Attention (GCA), which incorporates explicit 3D-aware priors for enhanced information fusion between image features and triplane tokens. Specifically, unlike the vanilla attention used in previous LRM works hong2023lrm ; wang2023pflrm ; li2023instant3d that accesses global features from all views, our method constructs an efficient attention mechanism. This mechanism ensures that each triplane token correlates with only a limited number of image feature tokens, improving computational efficiency and relevance.
Our intuition is that each token in a triplane corresponds to a ray in the 3D space passing through the token, which can be projected into each view (as illustrated in Fig. 4). In other words, each token is related to a projected ray within a view rather than the entire view. Consequently, when performing cross-attention between each triplane token and the image features, we only need to correlate the triplane token with the feature tokens intersecting each projected ray. In prior works, dense cross-attention operates on $n\times h\times w$ feature tokens, where $n$ is the number of views, and $h$ and $w$ are the height and width of the feature map. In contrast, our geometry-aware attention operates on only $n\times V$ feature tokens, where $V$ is the size of the geometry volume. This significantly reduces the computational complexity and focuses attention on more relevant features.
The key to geometry-aware attention is extracting the feature tokens that intersect with each projected ray. By retaining the grid volume as mentioned, we can leverage Eq. 1 to project the relevant volume vertices into each view and obtain the corresponding 2D locations. Using these projected 2D locations, we can then access the related feature tokens for cross-attention. This approach ensures that the attention mechanism focuses on the most pertinent features, enhancing efficiency and relevance in the information fusion process.
[…]
$Q=W^{Q}t^{s},K=W^{K}\left[f_{e,l}^{s}\right]_{e\in\mathcal{E},l\in\mathcal{L}},V=W^{V}\left[f_{e,l}^{s}\right]_{e\in\mathcal{E},l\in\mathcal{L}},$ (3)
where $\left[f_{e,l}^{s}\right]$ denote the features of the related projected feature tokens, $W^{Q}$, $W^{K}$ and $W^{V}$ are learnable matrices that project the inputs to query, key, and value, respectively.
To preserve global information interaction among different triplane tokens, we only replace half of the original transformer blocks with GCA, as it helps the model to preserve more details.
Figure 5: Illustration of Geometry-aware Cross Attention block (top) and Positional Embeddings (bottom).
3.3 Training Objectives
After we have the corresponding triplane $P$, we adopt differentiable NeRF renderer $\tau_{N}$ to render images and masks as $\{I^{pred},m^{pred}\}=\tau_{N}(E,K|P)$. The predicted novel views are supervised via differentiable rendering by other sampled views. We adopt MSE loss, binary mask loss and LPIPS loss zhang2018unreasonable . The training loss can be formulated as
$\mathcal{L}_{\text{NeRF }}$ $=\lambda_{\text{mse }}\sum_{i}\left\|I_{i}^{\text{pred }}-I_{i}^{\text{gt }}\right\|_{2}+\lambda_{\text{lpips }}\sum_{i}\mathcal{L}_{\text{lpips }}\left(I_{i}^{\text{pred }},I_{i}^{\text{gt }}\right)$ (4)
$+\lambda_{\text{mask }}\sum_{i}\left(m^{gt}\log m^{\text{pred }}+\left(1-m^{gt}\right)\log\left(1-m^{\text{pred }}\right).\right.$
4 Implementation Details and Datasets
Network Architecture. We train two variants of M-LRM, and here we provide details of our base variant. We refer to supplemental materials for more information on tiny variants.
For the image encoder, we utilize DINOv2-base with a patch size of 14. The transformer decoder consists of 12 layers with a hidden dimension of 768. We employ multi-head attention with 8 heads, and each head has a hidden dimension of 64. The learnable triplane tokens are initialized with a resolution of $64\times 64$ and 768 channels. To upsample the triplane, we use a transpose convolution layer with a kernel size of 2 and a stride of 2, resulting in a resolution of 128x128 and 40 channels. The concatenated triplane features are then decoded using a 4-layer MLP into density and RGB values. We sample 192 points per ray.
We construct the spatial index volume and cost volume with a resolution of $64^{3}$. After projecting the image features into spatial volumes, we aggregate them using a 5-layer 3D convolution networkwith channel sizes of [64, 128, 128, 128, 64]. Additionally, we derive additional triplane features from the cost volume using three convolution layers with a kernel size of 3 and a stride of 1.
Other Implementation Details. M-LRM is trained on 32 NVIDIA H800 GPU ( 80GB) for 3 days. The batch size for each GPU was set to 1, and gradient accumulation was employed for 8 steps, resulting in a total batch size of 512. At each iteration, we randomly sample 10 views among the total 24 images, where 6 of them are used as input conditional images, and the other 4 for supervision. AdamW optimizer is used with $\beta_{1},\beta_{2}$ set to $0.9,0.95$ respectively, and weight decay is set to 0.05. We increase the learning rate from 4e-10 to 4e-4 for the first epoch. After warmup, we adopt cosine annealing strategy for the learning rate and the learning rate decay drops to 0 at the end of training. The coefficients for different loss components are $\lambda_{mse}=1.0$, $\lambda_{L1}=0.2$, $\lambda_{mask}=0.1$ and $\lambda_{LPIPS}=1.0$. Following LGM tang2024lgm , we adopt camera augmentation and image grid distortion as data augmentation.
Figure 6: Qualitative result of sparse-view reconstruction. We compare our generated results with two baseline methods. Given the same input conditional views, our tiny model can generate results of higher fidelity and geometries that are more reasonable. Our base variant uses two additional views as input (the rightmost two images of the input views).
Dataset. We use the Objaverse dataset deitke2023objaverse . We filter out 190K objects among the 800K 3D models in the dataset. We use Blender to render 32 views per object. The 3D models are scaled to fit into the cubic bounding box, whose length is 1 and center is $(0,0,0)$. We set the field-of-view to $30^{\circ}$ and the camera distance is calculated accordingly to ensure the object is well fitted at the center of the view. The first 8 views are sampled uniformly around the object with $0^{\circ}$ elevation, the second 8 views are sampled uniformly with $20^{\circ}$ elevation while the third 8 views are sampled with $-20^{\circ}$ elevation. The last 8 views are sampled randomly on the surface of the sphere. All cameras face towards $(0,0,0)$.
5 Experiments
Figure 7: Visualization result of single view to 3D generation. We compare our method with Instant3D li2023instant3d , LGM tang2024lgm and TripoSR TripoSR2024 . The results of Instant3D suffer from blurry texture and severe geometry failure, while LGM tang2024lgm produces results with unreasonable color at the unseen region of the object. TripoSR can also generate visual results of high quality, but there remain many geometry artifacts. Our overall result is of better visual quality and geometry quality.
5.1 Multi-view Reconstruction
Evaluation Setting.
We compare our methods with Instant3D li2023instant3d and LGM tang2024lgm on Google Scanned Objects (GSO) downs2022google . Since the code and the model of Instant3D li2023instant3d are not available, we reproduce the model according to their paper. To fairly compare with Instant3D, we also train a tiny variant of our model that uses the same data as our reproduced Instant3D. We use the original implementation of LGM for the sparse view reconstruction task. Following existing work, we report PSNR, SSIM, and LPIPS to evaluate the visual quality of reconstruction and Chamfer distance (CD), F-Score(FS) for measuring the quality of the reconstructed mesh. We uniformly sample 16K points for measuring CD. We use two different threshold t to calculate FS for evaluation.
Results.
As shown in Tab. 1, our proposed M-LRM outperforms all baselines by a large margin. Even though our Base version uses the most views compared to previous methods, our tiny version model still demonstrates the superiority of our proposed method. Compared to Instant3D, the tiny M-LRM converges much faster, and can generate images of higher fidelity. As for the LGM that uses the same number of input views, though our tiny model uses much less data than LGM, we can still generate images of higher quality, i.e. higher PSNR. However, the lack of data makes the model fail to generalize to the evaluation dataset, leading to higher LPIPS. Fig. 6 shows some visualization results of sparse-view reconstruction on the GSO dataset. Our method can not only generate novel views of better visual quality with no floaters, but also is able to extract meshes that are more plausible.
Table 1: Comparison with state-of-the-art image-to-3D counterparts on the Google Scanned Objects (GSO) dataset. $\uparrow$ means the higher the better, and $\downarrow$ indicates the lower the better.
[…]
Instant3D li2023instant3d 4 256 22.18 0.958 0.142 0.0069 0.3357 0.7739
LGM tang2024lgm 4 512 21.74 0.967 0.110 0.0034 0.7108 0.9273
Ours (tiny) 4 256 24.35 0.961 0.142 0.0068 0.3357 0.771
Ours (Base) 6 512 25.23 0.977 0.096 0.0011 0.8725 0.9695
5.2 Single Image to 3D Generation
Evaluation Setting.
For single image to 3D generation, we adopt Zero123++ shi2023zero123++ to generate multiple views as conditional images. For our tiny variant and our reproduced Instant3D that takes 4 views as input, we select the first, third, fifth, and sixth images generated by Zero123++. As our model is trained with white background, while zero123++ generates multi-view images with gray background, we opt to use a refined version of zero123++ xu2024instantmesh that generates multi-view images with white background. We compare our methods with Instant3D li2023instant3d , LGM tang2024lgm , and TripoSR TripoSR2024 . For all methods, we render images at $512\times 512$ resolution.
Results.
We compare the qualitative results of our generated results with those of baseline methods. As shown in Fig. 7, we use five examples to explain our qualitative experiment. The generation outcomes of Instant3D exhibit deficiencies in texture clarity and significant faults (Cases 2 and 4) in geometric representation (cases 2, 3). Similarly, the results produced by LGM tang2024lgm display unreasonable texture color in the unobserved regions of the object (cases 1, 4, 5). Besides, the meshes produced by LGM contain some geometry artifacts. This is because of the complex pipeline for extracting mesh from 3D Gaussian representation. Although TripoSR is capable of generating visually appealing results, it still exhibits numerous imperfections in geometric accuracy (cases 1, 5) due to the lack of explicit 3D spatial modeling. In contrast, as our proposed method is geometry-aware, it yields superior visual quality and improved geometric fidelity.
5.3 Ablation Study
Table 2: The ablation study of each component.
[…]
24.501 0.9746 0.1364
24.923 0.9750 0.1322
25.116 0.9765 0.1269
26.219 0.9756 0.1201
25.254 0.9772 0.1246
26.966 0.9781 0.1067
27.581 0.9803 0.0986
We conduct ablation studies to demonstrate the effectiveness of our model components and architectural designs. We train these model variants with only Objaverse-LVIS subset. As shown in Tab. 2, GCA means using our proposed Multi-view Cross Attention module, GaPE means using Geometry-aware Positional Encoding, w/ Tr. Block means adding a transformer decoder block after our MCA block, and High Res. means using 64 as the resolution of the initial triplane and the geometry volume, as opposed to the original 32. When High Res. is adopted, we sample 192 points per ray in volumetric rendering instead of 96 originally.
The results of the first two rows and the last two rows demonstrate the effectiveness of our proposed GCA and GaPE. As shown in the table, adding a transformer decoder block after GCA block improves the model’s performance, indicating that preserving global attention is helpful. We also show that using High Res. setting can enhance the model’s capability by a large margin.
6 Conclusion and Limitation
Conclusion.
In this work, we introduce the Multi-view Large Reconstruction Model (M-LRM), which is designed to efficiently reconstruct high-quality 3D shapes from multiple views in a 3D-aware manner. Our proposed approach incorporates a multi-view consistent cross-attention scheme, enabling M-LRM to accurately extract relevant information from the input images. Additionally, we leverage the 3D priors of the input multi-view images to initialize the tri-plane tokens. By doing so, our model surpasses the capabilities of instant3D and produces 3D shapes with exceptional fidelity. Experimental results demonstrate that our model substantially improves performance and exhibits faster training convergence compared to instant3D and other LRM-series work.
Limitations and Future Work.
However, we notice that there are still some limitations in our framework. (i) Currently, our model can only accept a fixed number of views as input. However, our model can be further extended to support arbitrary input views, which will be left as future work. (ii) Although our model is trained with augmented data, its performance would still inevitably be affected by inconsistent multi-view input when doing single-view to 3D generation tasks. However, we believe that this problem can be solved as the multi-view generation models advance.